Karrie Co. has outstanding a 7%, 10-year bond with a $100,000 face amount. The bond was originally sold to yield 6% annual interest. Karrie uses the effective-interest-rate method to amortize bond premium. On June 30, year 1, the carrying amount of the outstanding bond was $105,000. What amount of unamortized premium on the bond should Karrie report in its June 30, year 2 balance sheet?Karrie Co. has outstanding a 7%, 10-year bond with a $100,000 face amount. The bond was originally sold to yield 6% annual interest. Karrie uses the effective-interest-rate method to amortize bond premium. On June 30, year 1, the carrying amount of the outstanding bond was $105,000. What amount of unamortized premium on the bond should Karrie report in its June 30, year 2 balance sheet?
Answer #1
Calculation of issue price of bond is shown below
Price of bond = Present value of coupon payment + Present value of face value
In this case we are assuming the interest is paid annually by company.
Year Cash flow Discount factor @ 6% Present value
1 $7,000 0.943396226 $6,603.77
2 $7,000 0.88999644 $6,229.98
3 $7,000 0.839619283 $5,877.33
4 $7,000 0.792093663 $5,544.66
5 $7,000 0.747258173 $5,230.81
6 $7,000 0.70496054 $4,934.72
7 $7,000 0.665057114 $4,655.40
8 $7,000 0.627412371 $4,391.89
9 $7,000 0.591898464 $4,143.29
10 $107,000 0.558394777 $59,748.24
Issue price $107,360.09
Coupon amount = 100000*7% = $7,000
At end of year 10, company would pay the face value of $100,000 and thus total cash flow would be $107,000
Thus, issue price of bond is $107,360.
Calculation of total premium on bond
Premium on bond = $107,360-$100,000
Premium on bond $7,360
Calculation of premium amortized till 30th June, 1st year
Premium amortized = $107,360-$105,000
Premium amortized $2,360
Thus, unamortized premium = (7360-2360)
Unamortized premium $5,000
Thus, unamortized premium is $5,000.